# Statistical noise enhances quantumness benefits in spin-network quantum reservoir computing

### v1

@@ III Results and Discussion
We now turn to the effect of Gaussian noise resulting from a finite number of measurements at the output layer. In Fig. 4, we compare the case without statistical noise with that subject to a statistical noise level of $\sigma=0.001$, which corresponds to a million measurements. We look at conditions corresponding to both an entanglement advantage (high input frequency in the presence of finite dissipation) and an entanglement disadvantage (low input frequency in the presence of finite dissipation) [36].
Statistical noise invariably reduces performance, albeit to greatly varying extents. The low-frequency regime appears to generally sustain a greater performance reduction, both in the presence and absence of dissipation. The high-frequency regime, on the other hand, exhibits a performance loss only at the spike at low but non-zero interaction strength. This spike is present only in the absence of statistical noise, and is more pronounced in the presence of dissipation, as reported in Ref. 36. By considering only the case without statistical noise, one may be tempted to conclude that dissipation might offer a computational advantage in that regime. However, our results here, which show that such a performance improvement does not hold in the presence of statistical noise, corroborate the argument made in Ref. 48 respecting the necessity of accounting for statistical noise resulting from a finite number of measurements.
We now proceed to intensify the effect of statistical noise (which amounts to making fewer measurements), first in the high-frequency regime that decidedly benefits from quantum entanglement. Fig. 5 shows, for an input frequency of $f=5$ and a dissipation strength of $\Gamma=0.01$, the total memory capacity for a wide range of statistical noise levels. Our most intense statistical noise level is $\sigma=0.3$, which corresponds to roughly 10 measurements. The memory capacity is invariably reduced by increasing the statistical noise, but the amount of performance loss is dependent on the amount of entanglement and coherence in the reservoir.
This observation raises an important question: what happens to the entanglement advantage as the rising statistical noise level causes the performance to drop? Our results in the left panel of Fig. 5 indicate that the entanglement advantage very much persists, notwithstanding the overall reduction in memory capacity. We also found it interesting to look at how performance behaves as a function of coherence rather than entanglement. These results are shown in the right panel of the same figure, and they tell a similar story, indicating that both entanglement and coherence can function as measures of quantumness whose presence may confer a computational advantage on reservoir computers. Specifically, there is a peak for both entanglement and coherence, but the peak is more pronounced as a function of coherence; this may be attributable to the continued growing of coherence in the localized regime. There seems to be an excessive amount of quantumness, as measured by either entanglement or coherence, that does not grant benefit. But since coherence continues to grow in the many-body localized phase while entanglement dies, as may be seen in Fig. 3, the performance peak is more defined with coherence on the x-axis.
@@
The behavior above evinces the possibility of turning an unfavorable trend into an unambiguously positive one. To further investigate this effect, we turn to the case of the intermediate input frequency of $f=1$, whose results we present in Fig. 7. Remarkably, the quantum indifference (or slight disadvantage) observable without accounting for statistical noise does indeed turn into a distinctly favorable trend. At this level of statistical noise, reservoirs with finite quantum entanglement and coherence do perform better than the unentangled, incoherent reservoir. However, performance saturation appears more distinctly in the case of quantum entanglement on the x axis, whereas there seems to be rather a decline in the case of coherence, similar to the high-frequency case.
IV Conclusions and Outlook
@@
We studied the effect of statistical noise on the performance of a spin network QRC system in linear memory tasks, and its implications with respect to the frequency of the input signal, the presence of dissipation in the system, and quantumness as measured by quantum entanglement and coherence.
@@
A more realistic model of statistical noise provides protection against possible spurious observations of performance enhancements due to, for instance, dissipation at low interaction strength. This emphasizes the importance of accounting for the reality of a finite number of measurements when employing real machines for QRC.
@@
In general, performance is degraded by statistical noise resulting from a finite number of measurements. However, substantially quantum reservoirs are found to be more resistant to the effect of statistical noise than their non-quantum counterparts. As a result, it appears that, under the constraint of having to take a small to moderate number of measurements, we observe a quantum benefit in certain regimes in which none was observed in the absence of statistical noise.
This quantumness was measured here by means of quantum entanglement and coherence. A moderate amount of either measure seems optimal under the conditions established as conducive to quantumness benefit, notwithstanding the continued growth of coherence into the many-body localized phase.
@@
Regardless of our choice of quantumness measure, we find that, within the limitation discussed above, statistical noise may elevate the system from a regime that does not benefit from quantumness to a regime that does. This suggests that constraints imposed by real implementations of QRC may, rather surprisingly, serve rather than obstruct the quest for the quantum advantage.
V Acknowledgments
This work was supported by the National Research Council through its Applied Quantum Computing Challenge Program, the Natural Sciences and Engineering Research Council (NSERC) of Canada through its NSERC Discovery Grant Program, the Alberta Major Innovation Fund, and Quantum City. We would also like to thank Khabat Heshami and Hadi Zadeh-Haghighi for the useful discussions and feedback.

### v2

@@ 5 Results and Discussion
We now turn to the effect of Gaussian noise resulting from a finite number of measurements at the output layer. We look at conditions corresponding to both an entanglement advantage (high input frequency in the presence of finite dissipation) and an entanglement disadvantage (low input frequency in the presence of finite dissipation), as detailed in Ref. kora_2024. As noted in Sec. 4B, our sequences are not i.i.d., and memory capacities are much higher at lower frequencies due to the slower variation of the input signal.
In Fig. 4, we intensify the effect of statistical noise (which amounts to making fewer measurements) in the high-frequency regime that decidedly performs better in the presence of quantum entanglement and coherence, as can be seen by comparing the performance of the leftmost points compared to the rest. This high frequency is chosen to be $f=5$ at a dissipation strength of $\Gamma=0.01$. We present the total memory capacity for a wide range of statistical noise levels. Our most intense statistical noise level is $\sigma=0.1$. The memory capacity is invariably reduced by increasing the statistical noise, but the amount of performance loss is dependent on the amount of entanglement and coherence in the reservoir.
This observation raises an important question: what happens to the entanglement advantage as the rising statistical noise level causes the performance to drop? Our results in the left panel of Fig. 4 indicate that the entanglement advantage very much persists, notwithstanding the overall reduction in memory capacity. We also found it interesting to look at how performance behaves as a function of coherence rather than entanglement. These results are shown in the right panel of the same figure, and they tell a similar story, indicating that both entanglement and coherence can function as measures of quantumness whose presence may correspond to a computational advantage on reservoir computers. Specifically, there is a peak for both entanglement and coherence, but the peak is more pronounced as a function of coherence; this may be attributable to the continued growing of coherence in the localized regime. There seems to be an excessive amount of quantumness, as measured by either entanglement or coherence, that does not correspond to much benefit. But since coherence continues to grow in the many-body localized phase while entanglement dies, as may be seen in Fig. 3, the performance peak is more defined with coherence on the x-axis.
@@
The behavior above evinces the possibility of turning an unfavorable trend into an unambiguously positive one. To further investigate this effect, we turn to the case of the intermediate input frequency of $f=1$, whose results we present in Fig. 6. Remarkably, the quantum indifference (or slight disadvantage) observable without accounting for statistical noise does indeed turn into a distinctly favorable trend. At this level of statistical noise, reservoirs with finite quantum entanglement and coherence do perform better than the unentangled, incoherent reservoir. However, performance saturation appears more distinctly in the case of quantum entanglement on the x axis, whereas there seems to be rather a decline in the case of coherence, similar to the high-frequency case. Thus, Fig. 6 presents a phenomenon where statistical noise causes a decrease in overall performance but the emergence of a relative advantage for reservoirs with more quantumness.
We make one last point by looking at memory capacity as a function of interaction strength in Fig. 7, both in the presence and the absence of noise ($\sigma=0.001$, which corresponds to a million measurements). As established, statistical noise invariably reduces performance, albeit to greatly varying extents. The low-frequency regime appears to generally sustain a greater performance reduction, both in the presence and absence of dissipation. The high-frequency regime, on the other hand, exhibits a performance loss only at the spike at low but non-zero interaction strength. This spike is present only in the absence of statistical noise, and is more pronounced in the presence of dissipation, as reported in Ref. kora_2024. By considering only the case without statistical noise, one may be tempted to conclude that dissipation might offer a computational advantage in that regime. However, our results here, which show that such a performance improvement does not hold in the presence of statistical noise, corroborate the argument made in Ref. palacios_2024 respecting the necessity of accounting for statistical noise resulting from a finite number of measurements.
6 Conclusions and Outlook
@@
Our central objective in this work was to investigate the behavior of our spin-network QRC system within the practical constraints of having to take a limited number of measurements. We studied the effect of the resulting statistical noise on the performance of our system in linear memory tasks, and its implications with respect to the frequency of the input signal, the presence of dissipation in the system, and quantumness as measured by quantum entanglement and coherence. Our chief conclusion is that, while the presence of statistical noise degrades performance on the whole, reservoirs that exhibit quantum entanglement and coherence are more resistant to this degradation that inevitably arises in implementations on present-day machines.
@@
A more realistic model of statistical noise also provides protection against possible spurious observations of performance enhancements due to, for instance, dissipation at low interaction strength. This emphasizes the importance of accounting for the reality of a finite number of measurements when employing real machines for QRC.
@@
In general, performance in general declines when the system is afflicted with the statistical noise resulting from a finite number of measurements. However, substantially quantum reservoirs are found to be more resistant to the effect of statistical noise than their non-quantum counterparts. As a result, it appears that, under the constraint of having to take a small to moderate number of measurements, we observe an entanglement advantage in certain regimes in which none was observed in the absence of statistical noise.
It is worth clarifying our use of the term ”entanglement advantage”, which in this study refers to the performance comparison between highly entangled states and low-entanglement states within quantum reservoirs, rather than an advantage over classical algorithms. It describes a phenomenon in which there is an observed improvement in QRC performance when there is non-zero entanglement. It does not, however, refer to a monotonic relationship where entanglement is the sole determiner of performance. For example, it is clear that stronger coupling between qubits can by itself influence reservoir memory by altering dynamical timescales; however, the observed trends across different interaction and dissipation regimes do not suggest that coupling is solely responsible either. For instance, it is clear from the right panels of Fig. 7 that high interaction strength leads to a decline in total memory capacity. Fig. 4 also shows that performance is not monotonically related to coherence and entanglement either, but that entangled and coherent reservoirs can be more resilient to statistical noise. The relative extents to which coupling and quantumness respectively correlate with performance is beyond the scope of the present study, but it is an interesting question for future investigation
It is also worth taking a moment to emphasize the distinction between our demonstrated quantum enhancement in noise tolerance on one hand, and the quantum advantage on the other. The latter refers to a computational advantage derived from quantum effects which cannot be replicated in a classical system, and its thorough demonstration is an ongoing quest. While our results here provide important considerations for that quest, they must be interpreted as evidence of entanglement-enhanced statistical noise resilience in quantum reservoir computers rather than of absolute quantum computational supremacy over classical reservoir computers.
We must stress again that our use of terms like “entanglement advantage” refers to regimes where entangled reservoirs enjoy a computational advantage relative to their non-entangled version, the latter still being a quantum system albeit in its classical limit. Such a quantum reservoir in its classical limit is still afflicted with the inevitability of statistical noise, and so incorporating it into the analysis is important for investigating the degree to which more entangled reservoirs may perform better than those with little to no quantum entanglement.
Our observed resilience to statistical noise is evocative of a similar phenomenon in continuous-variable QRC, where the quantum resource of squeezing was found to enhance the robustness of the reservoir to readout noise, giving rise to performance enhancement in linear and non-linear memory tasks garcia_2024.
@@
It is also worth reflecting on our reliance on primarily entanglement (and secondarily coherence) to quantify quantumness within our reservoirs. It may be of great interest to examine how this reliance may be affected by the existence of classically simulable operations with a high degree of entanglement, such as stabilizer states and their associated Clifford circuits gu_2025. A more nuanced measure of quantumness in light of this is the so-called magic, or nonstabilizerness, which may have implications for the possibility of reproducing the observed performance classically maronese_2025. A future study could employ magic quantifiers, such as the stabilizer 2-Rényi entropy leone_2022, to systematically analyze its presence and impact within our spin-network QRC scheme and how non-stabilizer operations might correlate with enhanced resilience to statistical noise.
In short, our results suggest that, within the limitations discussed above, the presence of statistical noise, while on the whole detrimental to performance, may nevertheless elevate the system from a regime that does not benefit from quantumness, as measured by entanglement and coherence, to a regime that does. This suggests that constraints imposed by real implementations of QRC may, rather surprisingly, serve rather than obstruct the search for QRC systems in which the presence of quantumness corresponds to higher performance.
Author Contributions
YK: conceptualization, methodology, investigation, computation, code development, manuscript writing; CS: conceptualization, supervision, validation, manuscript review and editing, funding acquisition. All authors have read and agreed to the final version of the manuscript.
Funding
This work was supported by the National Research Council through its Applied Quantum Computing Challenge Program, the Natural Sciences and Engineering Research Council (NSERC) of Canada through its NSERC Discovery Grant Program, the Alberta Major Innovation Fund, and Quantum City.
Acknowledgments
We would like to thank Khabat Heshami and Hadi Zadeh-Haghighi for the useful discussions and feedback.
Data Availability Statement
The datasets generated and analyzed during the current study are available from the corresponding author upon reasonable request.